OpenStudy (anonymous):

Predict the y-value for x = 6. X: 1 2 3 4 Y: 8.5 11 13.5 16 19.5 21 23.5 26

OpenStudy (anonymous):

21

OpenStudy (anonymous):

16 - 13.5 = 2.5 13.5 - 11 = 2.5 11- 8.5 = 2.5 See a pattern?

OpenStudy (anonymous):

That would be 21 because you are adding 2.5 basically to each y variable.

OpenStudy (anonymous):

THANKS GUYS

OpenStudy (anonymous):

No prob